OpenStudy (anonymous):

im really confused... In a 45°- 45°- 90° right triangle, one of the legs is 6 inches. How long are the other two sides?

OpenStudy (anonymous):

does it say if it is the hypotenuse?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can use the trig equations for right angled triangles do you know them?

OpenStudy (anonymous):

i do not believe so

OpenStudy (jamesj):

Well, this is a right angled triangle, so what do you know about that hypotenuse? Call it H. And call the other two sides A and B. What do you know about A, B and H?

OpenStudy (jamesj):

Most basic relationship for right-angled triangles: Pythagorus' Theorem: the square of the two right-angled sides equals the square of the hypotenuse I.e., A^2 + B^2 = H^2 Yes?

OpenStudy (jamesj):

still there?

OpenStudy (jamesj):

Now in addition, as the triangle is isosceles, the two side right angled sides, A and B are the same length.

OpenStudy (jamesj):

Are you following?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

So A^2 + B^2 = H^2 implies that A^2 + A^2 = 6^2 I.e., 2 A^2 = 36 or A^2 = 18 Now what then is A?

OpenStudy (anonymous):

since the two angles are both 45° and the other angle is 90° .. the legs are the same - which means both legs are 6 inches . you then you A2+B2=C2 ....

OpenStudy (anonymous):

i asked my certified math teacher :)

OpenStudy (jamesj):

Is a "leg" necessarily a right-angled side? Or is the hypotenuse? Havoc said 6 was the length of the hypotenuse.

OpenStudy (anonymous):

in the question he said the "leg is 6 inches" so if the hypotenuse is 6 , he needs to be more specific -.-

OpenStudy (jamesj):

So Havoc, we have this: If the hypotenuse = 6, then the other two sides are the same length, A and 2A^2 = 6^2 = 36 i.e. A^2 = 18 Therefore A = ..... what? If one of the right-angled sides is 6, A = 6 and then H^2 = A^2 + A^2 = 6^2 + 6^2 = 36 + 36 = 72 Therefore H = ..... what? So you need to figure out what 'leg' means; I don't know. When you do, use one these two calculations above and finish it.